Big Bob Corporation has a present capital structure consisting of common stock (10 million shares) and...

  1. Big Bob Corporation has a present capital structure consisting of common stock (10 million shares) and debt ($140 million, 6% coupon rate). The company needs to raise $46 million and is undecided between two financing plans.  Plan A: Equity financing. Under this plan, an additional common stock will be sold at $15 per share.  Plan B: Debt financing. Under this plan, the firm will issue 10% coupon bonds.  At what level of operating income (EBIT) will the firm be indifferent between the two plans? Assume a 40% marginal tax rate.
Answer #1
Big Bob Corporation
At the indifferent EBIT , the EPS for both the options will be
the same.We assume the indifferent EBIT is x
Particulars Plan A Plan B
Beginning no of outstanding common stocks                    10,000,000                10,000,000
New Equity price /share                                   15                                -  
Amount of Equity Raised                    46,000,000                                -  
No of Additional shares to be issued =$46M/$15=                      3,066,667                                -  
a Total No of outstanding stocks after Capital Increase                    13,066,667                10,000,000
Debt Value before additional capital raising                  140,000,000              140,000,000
Annual Interest @6%                      8,400,000                  8,400,000
Additional Debt raised                                    -                  46,000,000
Additional Interest payable@10%                                    -                    4,600,000
b Total Interest Expense /year                      8,400,000                13,000,000
Now Indifferent EBIT x x
Less Interest                      8,400,000                13,000,000
EBT (x-8,400,000) (x-13,000,000)
Tax @40% (x-8,400,000)*0.4 (x-13,000,000)*0.4
After Tax income (x-8,400,000)*0.6 (x-13,000,000)*0.6
Commom Shares Outstanding                    13,066,667                10,000,000
EPS = (x-8,400,000)*0.6/13066667 (x-13,000,000)*0.6/10000000
Now for Indifferent EBIT , EPS will be same under both options
Therefore , (x-8,400,000)*0.6/13066667 =(x-13,000,000)*0.6/10000000
x=28,000,000
So, The Indifferent EBIT between two plans =$28,000,000